|-3x-5|>=2
The answer should have form b < x < c
or the form x < b or x > c
Is all my paper says to do.
I dont understand HOW they want me to put it in that form and do the problem.
The topic on the paper says, Absolute Value Inequalities.
I would greatly appreciate it if someone coul dhelp me figur eout how to do this, and show me how they had done it so i can learn how to do it myself.
Thanks,
KKG.
: |-3x-5|>=2
The answer should have form b < x < c
or the form x < b or x > c
Is all my paper says to do.
I dont understand HOW they want me to put it in that form and do the problem.
The topic on the paper says, Absolute Value Inequalities.
I would greatly appreciate it if someone coul dhelp me figur eout how to do this, and show me how they had done it so i can learn how to do it myself.
Thanks,
KKG.

Answer by Edwin McCravy(2087)

 

(-3x-5)>=2

--------------------------------------------
Learn rules for removing absolute values from
inequalities of these forms:


Rule 1:
-------------------------------------------
      
abs(EXPRESSION)<N, abs(EXPRESSION)<=N (where N is not negative)

Write as 

-N < EXPRESSION < N, -N <= EXPRESSION <= N, respectively.

then solve. [If N is negative, there is no solution,
and solution set = Ø)


------------------------------------------------------
------------------------------------------------------

Rule 2:
-------------------------------------------
abs(EXPRESSION)>N, abs(EXPRESSION)>=N (where N is not negative)

Write as 

EXPRESSION < -NOREXPRESSION > N, EXPRESSION <= -NOREXPRESSION >= N
respectively.

Then solve each part.  (If N is negative, solution set is 
"all real numbers".)
------------------------------------------

You need Rule 2.  I gave you Rule 1 also because you
will need it for other problems:

abs(-3x-5)>=2

EXPRESSION = -3x-5

N=2

So write:

-3x-5 <= -2OR-3x-5 >= 2

Solve each part:

Add 5 to both sides of both parts:

-3x <= 3OR-3x >= 7

Divide both sides of both parts by -3,
Notice that dividing an inequality by a NEGATIVE
NUMBER reverses the inequality sign:

(-3x)/(-3) >= 3/(-3)OR(-3x)/(-3) <= 7/3

x >= -1ORx <= -7/3

So we shade the number line:

1. to the right of -1
and
2. to the left of -7/3, which is the same as -21/3.


<=======@-----------@==================>
 -3       -2       -1        0        1            

The circles should be darkened since the inequalities
are all underlined:

Solution in interval notation is 

(-infinity,-7/3] U [-1,infinity)

Edwin